OpenStudy (anonymous):

The equation of a circle is (x + 3)2 + (y + 5)2 = 8. Determine the coordinates of the center of the circle.

OpenStudy (anonymous):

Center (-3, -5) Center (3, 5) Center (-5, -3) Center (5, 3)

OpenStudy (anonymous):

you know these... :P

OpenStudy (anonymous):

Actually i have no idea lol

OpenStudy (anonymous):

lol...(h, k) is the center...so (x +3) means h = ? (y + 5) means k = ? what is your thought?

OpenStudy (anonymous):

B :D

OpenStudy (jdoe0001):

ahemm

OpenStudy (jdoe0001):

@PaolaMinaj have you bothered taking a peek a the standard form for the circle equation?

OpenStudy (anonymous):

lol your sure (:

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

(x + 3)2 + (y + 5)2 = 8. because its +3 it means its going to be -3 same with the =5 its actually -5 so the center would actually be A (-3,-5)

OpenStudy (anonymous):

do you get it ?